# Are There Correlations in the HAWC and IceCube High Energy Skymaps Outside the Galactic Plane?

### v1

@@ Abstract
Previous studies have searched for correlations between a neutrino flux and a gamma ray flux when one or both of these messengers are produced by identified point sources [13, 14]. However, there are likely many neutrino point sources which have remained hidden. Due to the inherent difficulty of detecting neutrinos, the number of neutrinos detected from each of these sources is small. We only consider the most energetic neutrino candidate events, which carry a higher probability of being neutrinos of astrophysical origin, and utilize events that are reconstructed as track-like, the signature of muon neutrinos, for good pointing accuracy. For our analysis we consider these muon neutrino candidates from the Southern hemisphere, observed as “down-going” events by IceCube, with reconstructed energy $E\geq{\rm PeV}$. Assuming each high-energy neutrino candidate is a potential source, we determine if there is a statistical excess of high energy gamma rays observed by HAWC [15] from these regions, compared to a similar sample of random sky regions. This analysis could allow us to detect a correlation between neutrinos and gamma rays which, though insignificant in any single sky region, could be statistically significant when summed over all such regions. A similar search for correlations between the unresolved gamma-ray background and neutrino events was considered in the context of Fermi-LAT gamma-ray data in the $1-25~{}{\rm GeV}$ range [4]. Our analysis will be sensitive to correlations at higher energies. Although there are reasons to expect that high energy astrophysical neutrinos are accompanied by much lower energy photons (see [16]), it is important to test the possibility that higher energy photons from these sources also arrive at Earth.
@@
We use the publicly-available data from the HAWC 3HWC point source catalog to obtain the signal significance for point source detection in each pixel, using data from 1523 days of operation [9]. The signal significance is the square of a test statistic given by $TS=2\ln({\cal L}_{s+b}/{\cal L}_{b})$, where ${\cal L}_{b}$ is the likelihood of the HAWC data given the HAWC background model, and ${\cal L}_{s+b}$ is the likelihood of the best fit model of background plus a point source with energy spectrum $\propto E^{-2.5}$, with the normalization as a free parameter [^1] The signal significance is reported with a negative sign if the best fit signal model has a negative signal normalization.. Since the test statistic in each pixel is $\chi^{2}$-distributed with one degree of freedom [9], the sum of the test statistic in 139 pixels should be $\chi^{2}$-distributed with 139 degrees of freedom. To demonstrate this, we draw $10^{4}$ sets of 139 random sky locations between declinations 00 and $-30^{\circ}$ (excluding the Galactic plane) and histogram the summed test statistics. In Figure 2, we show the histogram of summed test statistics for the random sky regions (blue) and the result for the 139 sky locations containing PeV neutrino candidates (orange line). For comparison, we also plot a $\chi^{2}$-distribution with 139 degrees of freedom (red curve), and the $\chi^{2}$-distribution which is a best fit to the histogram (black curve). The best fit $\chi^{2}$-distribution is for 143 degrees of freedom, similar to our expectations for the behavior of the test statistic. [^2] It is not surprising that there is a slight deviation between the distribution of the summed test statistic and the expected distribution under the background-only hypothesis, since there may be point sources in this region (that is, between declinations of 00 and $-30^{\circ}$, excluding the Galactic plane). The green and yellow bands are the $68\%$ and $95\%$ containment bands of the best fit $\chi^{2}$-distribution. We see that the test statistic obtained from the sky locations containing the IceCube neutrino candidates fits well inside the $68\%$ containment band, indicating that there is no statistical preference for point sources within these regions.
@@
Note that the HAWC angular resolution varies between $0.1^{\circ}-1.0^{\circ}$, and its pointing uncertainty may be as large as $0.3^{\circ}$ in this region of the sky [9]. The solid angle encompassed by 139 sky regions of size $0.3^{\circ}$ is small compared to the solid angle of the total sky region from which they are drawn. However, if the neutrino threshold were to be lowered significantly, one would obtain so many neutrinos that the regions of sky (of $0.3^{\circ}$ radius) covering these neutrino candidates would be an ${\cal O}(1)$ fraction of the intersection of the HAWC and IceCube footprints, largely negating the validity of this type of analysis. One might ask, however, if there is a correlation between gamma rays seen by HAWC and the highest energy neutrino candidates seen by IceCube. We investigate this possibility by performing a similar analysis using only the 18 neutrino candidate events detected by IceCube with $E>3.2~{}{\rm PeV}$, which are also within the HAWC footprint but outside the mask of the Galactic plane. We present the corresponding skymap in Figure 3, and the distribution of summed test statistics in Figure 4. We again find no significant statistical preference for gamma-ray point sources in this combined region.

### v2

@@ Abstract
We have described a few processes in which both photons and neutrinos are produced, with related spectra. However, photons and neutrinos may be produced through multiple processes that are not directly related to each other. As one example, dark matter particles may annihilate into multiple Standard Model final states. The subsequent decays of these Standard Model particles produce photons and neutrinos, along with other stable particles. The relationship between the photon energy spectrum and the neutrino energy spectrum would then depend in detail not only on the dark matter mass, but also on the branching fractions to the various final states. As another example, a source of high-energy neutrinos may also produce other cosmic rays, such as high-energy electrons or positrons, which can produce gamma rays through inverse Compton scattering off background photons.
Moreover, even the rough energy scale of the gamma rays at production, in comparison to the neutrino energy scale, depends on the production mechanism. For example, if photons and neutrinos are produced by the decays of neutral and charged pions, respectively, then their energy scales will tend to be comparable. On the other hand, if a process such as dark matter annihilation produces neutrinos, then the photons produced as a result of electroweak bremsstrahlung will be suppressed by in energy by up to an order of magnitude [2].
Furthermore, the photon spectrum will generally be reprocessed by interactions, both at the source and along the line of sight to the detector. For example, the high-energy photons can generate electromagnetic cascades by scattering off intervening background light (including starlight, CMB photons, etc) to produce $e^{+}e^{-}$ pairs, which in turn produce gamma-rays by inverse Compton scattering off background light. As such, the photon spectrum may be degraded in energy in comparison to that of the neutrinos. But how much reprocessing occurs will also depend on distance from the source to the Earth, as well as the details of the environment of the source.
We thus see that, given a specific model for the processes which produce neutrinos and photons at the source, one can obtain a prediction for the gamma ray spectrum, which can be searched for in data. But if one is agnostic about the production mechanism, one more generally sees the possibility that objects which emit high-energy neutrinos are also sources of gamma rays, as well as other cosmic rays. But as neutrinos and gamma rays are the messengers which will both point back to the source, there may be a correlation between the photon and neutrino skymaps. It is this possibility which we will study.
@@
Previous studies have searched for correlations between a neutrino flux and a gamma ray flux when one or both of these messengers are produced by identified point sources [13, 14]. However, there are likely many neutrino point sources which have remained hidden. Due to the inherent difficulty of detecting neutrinos, the number of neutrinos detected from each of these sources is small. We only consider the most energetic neutrino candidate events, which carry a higher probability of being neutrinos of astrophysical origin, and utilize events that are reconstructed as track-like, the signature of muon neutrinos, for good pointing accuracy. For our analysis we consider these muon neutrino candidates from the Southern hemisphere, observed as “down-going” events by IceCube, with reconstructed energy $E\geq{\rm PeV}$. Assuming each high-energy neutrino candidate is a potential source, we determine if there is a statistical excess of high energy gamma rays observed by HAWC [15] from these regions, compared to a similar sample of random sky regions. This analysis could allow us to detect a correlation between neutrinos and gamma rays which, though insignificant in any single sky region, could be statistically significant when summed over all such regions. A similar search for correlations between the unresolved gamma-ray background and neutrino events was considered in the context of Fermi-LAT gamma-ray data in the $1-25~{}{\rm GeV}$ range [4]. Our analysis will be sensitive to correlations at higher energies. Although there are reasons to expect that high energy astrophysical neutrinos are accompanied by much lower energy photons (see [16, 17]), it is important, for the reasons discussed above, to test the possibility that higher energy photons from these sources also arrive at Earth.
@@
We use the publicly-available data from the HAWC 3HWC point source catalog to obtain the signal significance for point source detection in each pixel, using data from 1523 days of operation [9]. The signal significance is the square of a test statistic given by $TS=2\ln({\cal L}_{s+b}/{\cal L}_{b})$, where ${\cal L}_{b}$ is the likelihood of the HAWC data given the HAWC background model [^1] For a point source search, the HAWC collaboration estimates background (dominated by misidentified hadronic cosmic rays) using the “direct integration” method detailed in Ref. [19]. The background rate in any region of sky is found by convolving an isotropic but time-dependent background rate with an acceptance factor which depends on the orientation with respect to the detector. Both of these factors are derived from HAWC data., and ${\cal L}_{s+b}$ is the likelihood of the best fit model of background plus a point source with energy spectrum $\propto E^{-2.5}$, with the normalization as a free parameter [^2] The signal significance is reported with a negative sign if the best fit signal model has a negative signal normalization.. Since the test statistic in each pixel is $\chi^{2}$-distributed with one degree of freedom [9], the sum of the test statistic in 139 pixels should be $\chi^{2}$-distributed with 139 degrees of freedom. To demonstrate this, we draw $10^{4}$ sets of 139 random sky locations between declinations 00 and $-30^{\circ}$ (excluding the Galactic plane) and histogram the summed test statistics. In Figure 2, we show the histogram of summed test statistics for the random sky regions (blue) and the result for the 139 sky locations containing PeV neutrino candidates (orange line). For comparison, we also plot a $\chi^{2}$-distribution with 139 degrees of freedom (red curve), and the $\chi^{2}$-distribution which is a best fit to the histogram (black curve). The best fit $\chi^{2}$-distribution is for 143 degrees of freedom, similar to our expectations for the behavior of the test statistic. [^3] It is not surprising that there is a slight deviation between the distribution of the summed test statistic and the expected distribution under the background-only hypothesis, since there may be point sources in this region (that is, between declinations of 00 and $-30^{\circ}$, excluding the Galactic plane). The green and yellow bands are the $68\%$ and $95\%$ containment bands of the best fit $\chi^{2}$-distribution. We see that the test statistic obtained from the sky locations containing the IceCube neutrino candidates fits well inside the $68\%$ containment band, indicating that there is no statistical preference for point sources within these regions.
@@
Note that the HAWC angular resolution varies between $0.1^{\circ}-1.0^{\circ}$, and its pointing uncertainty may be as large as $0.3^{\circ}$ in this region of the sky [9]. For the IC neutrino sample we use, in the energy range we consider, the median neutrino pointing uncertainty is $\sim 0.2^{\circ}$ [1]. Our method amounts to a search for excess gamma ray emission from the reconstructed directions of high-energy IC neutrino candidate events. This method would not be sensitive to correlations between the neutrino and gamma ray fluxes if the reconstructed neutrino direction is sufficiently far from the actual source. Since the pointing uncertainties are comparable, a small but non-negligible fraction of $\nu-/\gamma-$sources would evade this search in this manner.
The solid angle encompassed by 139 sky regions of size $0.3^{\circ}$ is small compared to the solid angle of the total sky region from which they are drawn. However, if the neutrino threshold were to be lowered significantly, one would obtain so many neutrinos that the regions of sky (of $0.3^{\circ}$ radius) covering these neutrino candidates would be an ${\cal O}(1)$ fraction of the intersection of the HAWC and IceCube footprints, largely negating the validity of this type of analysis. One might ask, however, if there is a correlation between gamma rays seen by HAWC and the highest energy neutrino candidates seen by IceCube. We investigate this possibility by performing a similar analysis using only the 18 neutrino candidate events detected by IceCube with $E>3.2~{}{\rm PeV}$, which are also within the HAWC footprint but outside the mask of the Galactic plane. We present the corresponding skymap in Figure 3, and the distribution of summed test statistics in Figure 4. We again find no significant statistical preference for gamma-ray point sources in this combined region.
@@
Not all sources of both gamma-rays and neutrinos are amenable to the search we have conducted. In particular, we have searched for spatial correlations between high energy neutrino candidates ($>{\rm PeV}$) seen by IceCube and lower energy photons seen by HAWC. Depending on the production mechanisms for these messengers, the distance to the source, and the environment between the source and Earth, even a source of both gamma rays and neutrinos may not produce detectable event rates within the available energy windows for these messengers. The limited publicly-available data thus limits the information one could obtain about the nature of the sources to which a search of this type is sensitive, beyond the important fact of their existence. When more data is publicly available, a future study involving spectral information from broader data sets would allow constraints on (or detections of) specific classes of sources and messenger production mechanisms.